OpenStudy (anonymous):

Help please :)? (-2/5)x-2=-22

OpenStudy (anonymous):

wut?

rishavraj (rishavraj):

(-2/5)x = -20 (2/5)x = 20....proceed

OpenStudy (anonymous):

(-2/5)x-2=-22 (-2/5)x=-22+2 (-2/5)x = -20 (-5/2)((-2/5)x=-20) x=50